a father is 30years iod and his son is 14 years; when in the past was the product of their age 80?
Answer by mananth(16946)

a father is 30years old and his son is 14 years; when in the past was the product of their age 80?
Let the product of their ages be 80, x years ago
(30-x)(14-x) =80
420 -30x-14x+x^2=80
x^2-44x+340=0
(x - 34)(x - 10)=0
x=34 or 10
x=34 is absurd
so x=10